Two angles with measures x2 + 5x + 100 and x2 + 9x - 40 are supplementary. Find the value of x and the measure of each angle

I am basically ripping my hair out trying to answer this and failing EPICLY! An answer ASAP would be AMAZING! I frikkin hate math! Makes me feel so stupid sometimes I swear!

Answer by mananth(16949)

x2 + 5x + 100 and x2 + 9x - 40
since they are supplementary sum = 180 degrees
x2 + 5x + 100 + x2 + 9x - 40=180
2x^2+14x+60=180
2x^2+14x-120=0
/2
x^2+7x-60=0
x^2+12x-5x-60=0
x(x+12)-5(x+12)=0
(x+12)(x-5)=0
x=5 OR -12
150 &30